What separation distance do two point charges of +1.0µC and −1.0µC exerts a force of attraction on each that is 440N? 


1
Expert's answer
2021-06-29T08:21:31-0400

According to Coulomb's Law,

F = "\\frac{k.q1.q2\ufeff}{r^2}"

Here,

F = 440 N

k = 9 x 109

q1 = 1 x 10-6 C

q2 = -1 x 10-6 C

Hence,

r2 = "\\frac{k.q1.q2\ufeff}{F}"

= (9 x 109 x 1 x 10-6 x 1 x 10-6) / 440

r2 = "\frac{9}{440000}"

Hence,

r = 0.0045 m

= 4.5 mm